Latron believes if he cuts back on his part-time job, he can spend some additional time
working out over the next 20 weeks and earn an athletic scholarship to the local college.
However, cutting back at work means he will forgo making $12 an hour at his uncle’s
shop. Assuming he now works only 10 hours a week with his reduced schedule instead of
the normal 22 hours a week, how much will his scholarship have to pay for him to break
even?
Can it be solved ASAP

Answer by ankor@dixie-net.com(22740)

Latron believes if he cuts back on his part-time job, he can spend some additional time working out over the next 20 weeks and earn an athletic scholarship to the local college.
However, cutting back at work means he will forgo making $12 an hour at his uncle’s shop.
Assuming he now works only 10 hours a week with his reduced schedule instead of
the normal 22 hours a week, how much will his scholarship have to pay for him to break even?
:
let s = amt of scholarship
s + 20(10*12) = 20(22*12)
s + 2400 = 5280
s = 5280 - 2400
s = $2880 scholarship to break even